# Topological Strings in SU(3) Gauge Theory at Finite Temperature

### v1

@@ Abstract
We investigate string configurations in the deconfined phase of SU(3) gauge theory, which arise from the spontaneous breaking of the $Z_{3}$ center symmetry. These configurations form at the junctions of domain walls of the theory. The complex phase of the Polyakov loop changes by multiples of $2\pi$ on large spatial loops around the string, rendering them topologically stable. Using the Monte Carlo simulations of the partition function, we compute the free energy associated with these configurations. The simulations are performed on lattices with spatial dimensions $N_{x,y}=60,N_{z}=4$, and temporal extent $N_{\tau}=2$. Our results show that the free energy of the $Z_{3}$-strings is dominated by the domain walls. Further near the transition point, thermal fluctuations cause the decay of domain walls as well as the $Z_{3}$ strings into confined-deconfined interfaces.
@@ I Introduction
The non-perturbative study of Quantum Chromodynamics (QCD), plays a crucial role in understanding matter under extreme conditions. In particular, it provides key insights into the phase diagram and the transition from the hadronic phase to the quark-gluon plasma (QGP). The thermodynamics of QCD remains an active field of research, driven by relativistic heavy-ion collision(HIC) experiments that probe different regions of the phase diagram. While the chiral and heavy-quark limits do not correspond to physical mass parameters, investigating these limits has significantly enhanced our understanding on the role and possible effects of chiral symmetry breaking and the confinement-deconfinement aspects in realistic QCD Satz:1982fu ; Kallman:1982yc ; Celik:1983nc ; DeGrand:1986tp ; Letessier:1994wm ; Kalashnikov:1996nf ; Bali:1993fb ; Caselle:1994df ; Belyaev:1991gh ; Lenz:1998eu ; Gazdzicki:1999ej ; Baldo:2003id ; Fiore:2004pg ; Nefediev:2009kn ; Greensite:2012dy ; Sakai:1997ia ; Borsanyi:2010cw ; APE:1987ehd ; Bali:2005fu ; Cornwall:1982zn ; Dubin:1994vn ; Petreczky:2012rq ; DeTar:2009ef ; Baker:2024peg ; Bicudo:2011hn .
@@
QCD in the heavy-quark limit, i.e., pure $SU(3)$ gauge theory, has been extensively studied using lattice simulations, which offer a powerful approach to explore the non-perturbative regime. In finite-temperature studies of this theorySvetitsky:1982gs ; Celik:1983wz ; Boyd:1996bx ; Iwasaki:1992ik ; Lucini:2005vg , it has been firmly established that the system undergoes a first-order confinement-deconfinement (CD) phase transition, at the critical temperature ($T_{c}$), between the hadronic phase and the QGP phase. The hadronic and the QGP phases can be interpreted in terms of glueballs and a thermalized medium of gluons, respectively. In pure $SU(N)$ gauge theories, the Polyakov loop expectation value($L$), which transforms non-trivially under the $Z(N)$ gauge transformations, serves as the order parameter for the CD phase transition. In the deconfined phase, the Polyakov loop, acquires a nonzero expectation value, i.e., $L\neq 0$, leading to spontaneous breaking of the $Z(N)$ symmetry. Consequently there are $N$ degenerate states. For $N\geq 3$, the Polyakov loop is complex valued and these states are characterised by complex phases of the Polyakov loop, given by,
$\theta_{k}={2\pi k\over N},\leavevmode\nobreak\ k=0,1,..,N-1.$ (1)
@@
In contrast, for $N=2$, the Polyakov loop is real-valued, and the states are distinguished by its sign. These vacua of discrete states give rise to static domain walls (interfaces) that interpolate between a pair of vacuum sectors, i.e., $\theta_{i}$’s Mermin:1979rmp .
@@
The domain walls in Pure $SU(3)$ gauge theory, have been studied using lattice simulations, specifically for $N_{\tau}=2$ Kajantie:1990bu ; Aoki:1993dm ; Holland:1999mx and for $N_{\tau}=4$ Huang:1990jf . A more recent study on the calculation of interface tension for $N_{\tau}\geq 2$ has been conducted using different numerical algorithms and employing the t’Hooft loop dual operator, with significant findings reported in deForcrand:2005pb ; deForcrand:2004jt ; deforcrand:2005 . Details on the $2+1D$ simulation of the $Z_{3}$ interfaces can be found in West:1996sb ; West:1996ej ; Korthals_Altes:1995 ; these studies provide a comprehensive numerical framework that captures the complex behaviour of these interfaces under various conditions. In earlier work, the interface tension was calculated within a perturbative framework for $SU(N)$ gauge theories at finite temperatures, and this was achieved by establishing a connection between the $Z(N)$ vacua and the instanton solutions of an effective action that incorporates both classical and quantum fluctuations Bhattacharya:1992qb ; Weiss:1980rj .
@@
For $SU(N)$ gauge theory in the deconfined phase, the interface tension, $\alpha_{ij}$, of the domain walls depends on the difference in the phases of the Polyakov loop, $|\theta_{i}-\theta_{j}|$, apart from temperature (T) deForcrand:2004jt ; Giovannangeli:2001bh . Further, $\alpha_{ij}+\alpha_{jk}\geq\alpha_{ik}$, thus the interaction between these domain walls is attractive. The Polyakov loop takes on different thermal expectation values on either side of the domain walls. Its magnitude decreases as one approaches the center from either side of the domain wall, where it reaches a non-zero minimum value (see Section II). As a result, in a configuration of multiple domain walls, the phase of the Polyakov loop remains well-defined in most of the physical space. This suggests the possibility of an intriguing configuration in the deconfined phase, at the intersection of all the $N$ vacua. The domain walls can be oriented such that, the topological considerations impose a line-like intersection Layek:2005fn ; Balachandran:2001qn , where the value of the Polyakov loop $L(\vec{x})$ becomes exactly zero. As $L(\vec{x})$ becomes zero, locally there is restoration of $Z_{N}$ symmetry and confining behaviour in an overall deconfined environment. Thus, at high temperatures, the vacuum structure is such that the multiple degenerate $Z(N)$ states meet to form a defect whose core resembles the confined phase. This picture has been further elucidated by studies that employ lattice simulations and effective models to explore nonperturbative features of the quark-gluon plasma. In particular, for $N=3$, string configurations were investigated Layek:2005fn using an effective potential for the Polyakov loop with $Z_{3}$-symmetry Pisarski:2000eq , providing valuable insights into how the order parameter deviates sharply from the value corresponding to the deconfined phase, due to the underlying topological structure. These results, which are highly dependent on the effective potential, can only be validated through first-principles calculations, such as lattice simulations of the thermal partition function.
@@
In this paper, we explore the topological string configurations in $SU(3)$ gauge theory using lattice simulations, investigating their intricate role in elucidating the non-perturbative phenomena inherent in QCD. The free energy for these configurations is computed by integrating the action difference between systems with and without the string, a standard technique in statistical physics. In our study we also compute the interface tension, which is in good agreement with previous results Kajantie:1990bu , reinforcing the consistency of our approach with earlier studies and highlighting the relevance of these methods in probing non-trivial topological configurations in gauge theories.
@@
The $Z_{3}$ topological string studied in this paper and flux-tubes in QCD are fundamentally different. In the case of the $Z_{3}$ string, it is a deconfined phase everywhere except at the core. The situation reverses for flux-tubes, i.e., the phase outside is confined and is deconfined inside. The QCD strings that emerge in the confining phase of quantum chromodynamics (QCD), where they manifest as flux tubes linking quark-antiquark pairs or forming glueball states within the framework of pure gluonic theory—a perspective that is continually refined by numerical and analytical studies Hansson:1982dv ; APE:1987ehd . Unlike the QCD strings, which are prone to breaking via the spontaneous creation of quark-antiquark pairs Detar:1998qa ; Bali:1992ab ; Nefediev:2009kn ; Johnson:2000qz , the strings we examine are topologically stable, which allows them to persist and interact in the deconfined phase.
@@
In this study, the effects of the dynamical quarks on the string configurations are neglected. However, previous research provides substantial evidence suggesting that quarks can influence such configurations. Both perturbative and mean-field studies indicate that the presence of quarks explicitly breaks the $Z_{3}$ symmetry Green:1983sd ; Hasenfratz:1983ce ; Satz:1985js ; Belyaev:1991np . The explicit breaking weakens the CD transition from a pure gauge first order transition to a crossover with decrease in the quark masses in the heavy quark regime Green:1983sd ; Hasenfratz:1983ce ; Satz:1985js . The explicit breaking lifts the degeneracy between the vacua. Since the three domain walls, in other words three vacua, join to form the string configuration, the lifting of degeneracy between the $Z_{3}$ vacua renders the string non-static. It will move towards the region of meta-stable states so as to reduce the free energy of the system Gupta:2011ag . In the presence of dynamical quarks, near $T_{c}$, only the $\theta=0$ state is observed in the deconfined phase. The states with $\theta=\pm 2\pi/3$ appear as metastable states only above a certain temperature $T_{m}$ Deka:2012 . This suggests that the string configuration will become unstable and ”melt” as the temperature drops below $T_{m}$. Analysis within the framework of the PNJL model at zero baryon chemical potential shows that if metastable states were to form in heavy-ion collision(HIC) experiments, they will survive until the temperature falls below $T_{m}$ Biswal:2019xju . This suggests that string configurations may potentially affect the dynamics of QGP above $T_{m}$ in HIC.
@@ II $Z_{3}$-string in pure $SU(3)$ gauge theory
In this section, we briefly discuss the $Z_{3}$ symmetry in pure $SU(3)$ gauge theory at finite temperature and how topological defects follow from the Polyakov loop effective potential. In pure $SU(3)$, the gauge fields $A_{\mu}$ at any point in space-time, are matrices in color space, i.e., $A_{\mu}=A^{a}_{\mu}T^{a}$, where $T^{a},a=1,2...8$, are generators of $SU(3)$ Georgi:1999wka . The partition function in the path-integral formulation is given by,
@@
Consequently, the effective potential has three degenerate minima, as illustrated in Fig.1. In the complex $L$-plane, the point $L=0$ corresponds to a saddle point of the effective potential, while the minima occur for $|L|\neq 0$ with the polar coordinate(phase) $\theta=0,2\pi/3$, and $2\pi/3$ Pisarski:2001pe ; Weiss:1980rj , which we denote by $L_{0},L_{1}$ and $L_{2}$ respectively. Given this, there are three possible domain walls, i.e., $L_{ij}$ interpolating $L_{i}$ and $L_{j}$. A static domain wall solution, on the yzyz-plane, is obtained by solving the following field equations,
${d^{2}l_{i}\over dx^{2}}={\partial V\over\partial l_{i}},\leavevmode\nobreak\ i=1,2,$ (6)
@@
$\lim_{x\to-\infty}L(x)=L_{i},\leavevmode\nobreak\ \lim_{x\to+\infty}L(x)=L_{j}.$
@@ III $SU(3)$ gauge theory on the lattice
We consider the standard Wilson action Boyd:1996bx , as the discrete formulation of the action, Eq.2 , on a four dimensional Eucledian lattce,
$S=\beta\sum_{n,\mu>\nu}Tr[1-\frac{1}{3}Re\leavevmode\nobreak\ U_{\mu\nu}],$ (7)
where $U_{\mu\nu}=U_{\mu}(n)U_{\nu}(n+\hat{\mu}){U^{\dagger}}_{\mu}(n+\hat{\nu}){U^{\dagger}_{\nu}(n)}$ is the standard Wilson plaquette with $\mu,\nu=1,2,3,4$, $n=(\vec{n},n_{4})$. $\hat{\mu}$ is a vector of length equal to the lattice spacing($a$) in the $\mu$-th direction. $U_{\mu}(n)=\exp[igaA_{\mu}(n)]$ is the link variable, resides on the link connecting the lattice points $n$ and $n+\hat{\mu}$. $\beta=6/g^{2}$ is the lattice gauge coupling that regulates the lattice constant ($a$) and the temperature scale through the $\beta$-function.
On the lattice, the trace of the product of links in the temporal direction gives the Polyakov loop operator,
$L(\vec{n})=\frac{1}{3}Tr\prod_{n_{4}=1}^{N_{\tau}}U_{4}(\vec{n},n_{4})$ (8)
For the realisation of the string configurations, we use the Polyakov loop as a function of two spatial directions (x,y), which is given by,
$L(n_{1},n_{2})=\frac{1}{N_{z}}\sum_{n_{3}=1}^{N_{z}}\frac{1}{3}Tr\prod_{n_{4}=1}^{N_{\tau}}U_{4}(n_{1},n_{2},n_{3},n_{4}),$ (9)
@@
To compute the free energy of strings, appropriate boundary conditions are imposed on the lattice to induce a specific string configuration, as discussed in the next section. Specifically, we evaluate the free energy difference between the configurations with and without strings. The free energy can be calculated from the partition function, F=TlnZF=-TlnZ, where $Z$ is the partition function and $T$ is the temperature, but in practice, the direct evaluation of the partition function in lattice simulations is avoided as it is difficult. Instead, we evaluate the derivative of the free energy $w.r.t.$ the lattice gauge coupling $\beta$, which is related to the action difference between simulations with and without the string, i.e.,
$\frac{\partial}{\partial\beta}\left(\frac{F}{T}\right)$ $=\frac{1}{\beta}\langle\Delta S\rangle,$ (10)
where $\Delta S=S_{1}-S_{0}$, and $S_{1}$ and $S_{0}$ are actions with and without string respectively. The free energy is obtained by integrating the expectation value,
$\frac{F}{T}{\Big{|}}_{\beta_{c}}^{\beta}$ $=$ $\int_{\beta_{c}}^{\beta}d\beta\leavevmode\nobreak\ \left<{1\over\beta}\Delta S\right$> (11)
$=$ $\int_{\beta_{c}}^{\beta}d\beta\leavevmode\nobreak\ \left<\Delta\left(\sum_{n,\mu>\nu}Tr[1-\frac{1}{3}Re\leavevmode\nobreak\ U_{\mu\nu}]\right)\right$>

### v2

@@ Abstract
We investigate string configurations in the deconfined phase of SU(3) gauge theory, which arise from the spontaneous breaking of the $Z_{3}$ center symmetry. These configurations form at the junctions of domain walls of the theory. The complex phase of the Polyakov loop changes by multiples of $2\pi$ on large spatial loops around the string, rendering them topologically stable. Using the Monte Carlo simulations of the partition function, we compute the free energy associated with these configurations. The simulations are performed on lattices with spatial dimensions $N_{x,y}=60,N_{z}=4$, and temporal extent $N_{\tau}=2,4$. Our results show that the free energy of the $Z_{3}$-strings is dominated by the domain walls. Further near the transition point, thermal fluctuations cause the decay of domain walls as well as the $Z_{3}$ strings into confined-deconfined interfaces.
@@ I Introduction
The non-perturbative study of Quantum Chromodynamics (QCD), plays a crucial role in understanding matter under extreme conditions. In particular, it provides key insights into the phase diagram and the transition from the hadronic phase to the quark-gluon plasma (QGP). The thermodynamics of QCD remains an active field of research, driven by relativistic heavy-ion collision(HIC) experiments that probe different regions of the phase diagram. While the chiral and heavy-quark limits do not correspond to physical mass parameters, investigating these limits has significantly enhanced our understanding on the role and possible effects of chiral symmetry breaking and the confinement-deconfinement aspects in realistic QCD [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
@@
QCD in the heavy-quark limit, i.e., pure $SU(3)$ gauge theory, has been extensively studied using lattice simulations, which offer a powerful approach to explore the non-perturbative regime. In finite-temperature studies of this theory[26, 27, 28, 30, 29], it has been firmly established that the system undergoes a first-order confinement-deconfinement (CD) phase transition, at the critical temperature ($T_{c}$), between the hadronic phase and the QGP phase. The hadronic and the QGP phases can be interpreted in terms of glueballs and a thermalized medium of gluons, respectively. In pure $SU(N)$ gauge theories, the Polyakov loop expectation value($L$), which transforms non-trivially under the $Z(N)$ gauge transformations, serves as the order parameter for the CD phase transition. In the deconfined phase, the Polyakov loop, acquires a nonzero expectation value, i.e., $L\neq 0$, leading to spontaneous breaking of the $Z(N)$ symmetry. Consequently there are $N$ degenerate states. For $N\geq 3$, the Polyakov loop is complex valued and these states are characterised by complex phases of the Polyakov loop, given by,
$\theta_{k}={2\pi k\over N},\penalty 10000\ k=0,1,..,N-1.$ (1)
@@
In contrast, for $N=2$, the Polyakov loop is real-valued, and the states are distinguished by its sign. These vacua of discrete states give rise to static domain walls (interfaces) that interpolate between a pair of vacuum sectors, i.e., $\theta_{i}$’s [31].
@@
The domain walls in Pure $SU(3)$ gauge theory, have been studied using lattice simulations, specifically for $N_{\tau}=2$ [32, 33, 34] and for $N_{\tau}=4$ [35]. A more recent study on the calculation of interface tension for $N_{\tau}\geq 2$ has been conducted using different numerical algorithms and employing the t’Hooft loop dual operator, with significant findings reported in [36, 38, 37, 39]. Details on the $2+1D$ simulation of the $Z_{3}$ interfaces can be found in [40, 41, 42]; these studies provide a comprehensive numerical framework that captures the complex behaviour of these interfaces under various conditions. In earlier work, the interface tension was calculated within a perturbative framework for $SU(N)$ gauge theories at finite temperatures, and this was achieved by establishing a connection between the $Z(N)$ vacua and the instanton solutions of an effective action that incorporates both classical and quantum fluctuations [45, 43].
@@
For $SU(N)$ gauge theory in the deconfined phase, the interface tension, $\alpha_{ij}$, of the domain walls depends on the difference in the phases of the Polyakov loop, $|\theta_{i}-\theta_{j}|$, apart from temperature (T) [38, 47]. Further, $\alpha_{ij}+\alpha_{jk}\geq\alpha_{ik}$, thus the interaction between these domain walls is attractive. The Polyakov loop takes on different thermal expectation values on either side of the domain walls. Its magnitude decreases as one approaches the center from either side of the domain wall, where it reaches a non-zero minimum value (see Section II). As a result, in a configuration of multiple domain walls, the phase of the Polyakov loop remains well-defined in most of the physical space. This suggests the possibility of an intriguing configuration in the deconfined phase, at the intersection of all the $N$ vacua. The domain walls can be oriented such that, the topological considerations impose a line-like intersection [48, 49], where the value of the Polyakov loop $L(\vec{x})$ becomes exactly zero. As $L(\vec{x})$ becomes zero, locally there is restoration of $Z_{N}$ symmetry and confining behaviour in an overall deconfined environment. Thus, at high temperatures, the vacuum structure is such that the multiple degenerate $Z(N)$ states meet to form a defect whose core resembles the confined phase. This picture has been further elucidated by studies that employ lattice simulations and effective models to explore nonperturbative features of the quark-gluon plasma. In particular, for $N=3$, string configurations were investigated [48] using an effective potential for the Polyakov loop with $Z_{3}$-symmetry [50], providing valuable insights into how the order parameter deviates sharply from the value corresponding to the deconfined phase, due to the underlying topological structure. These results, which are highly dependent on the effective potential, can only be validated through first-principles calculations, such as lattice simulations of the thermal partition function.
@@
In this paper, we explore the topological string configurations in $SU(3)$ gauge theory using lattice simulations, investigating their intricate role in elucidating the non-perturbative phenomena inherent in QCD. The free energy for these configurations is computed by integrating the action difference between systems with and without the string, a standard technique in statistical physics. In our study we also compute the interface tension, which is in good agreement with previous results [32], reinforcing the consistency of our approach with earlier studies and highlighting the relevance of these methods in probing non-trivial topological configurations in gauge theories.
@@
The $Z_{3}$ topological string studied in this paper and flux-tubes in QCD are fundamentally different. In the case of the $Z_{3}$ string, it is a deconfined phase everywhere except at the core. The situation reverses for flux-tubes, i.e., the phase outside is confined and is deconfined inside. The QCD strings that emerge in the confining phase of quantum chromodynamics (QCD), where they manifest as flux tubes linking quark-antiquark pairs or forming glueball states within the framework of pure gluonic theory—a perspective that is continually refined by numerical and analytical studies [46, 18]. Unlike the QCD strings, which are prone to breaking via the spontaneous creation of quark-antiquark pairs [61, 59, 14, 60], the strings we examine are topologically stable, which allows them to persist and interact in the deconfined phase.
@@
In this study, the effects of the dynamical quarks on the string configurations are neglected. However, previous research provides substantial evidence suggesting that quarks can influence such configurations. Both perturbative and mean-field studies indicate that the presence of quarks explicitly breaks the $Z_{3}$ symmetry [52, 53, 54, 55]. The explicit breaking weakens the CD transition from a pure gauge first order transition to a crossover with decrease in the quark masses in the heavy quark regime [52, 53, 54]. The explicit breaking lifts the degeneracy between the vacua. Since the three domain walls, in other words three vacua, join to form the string configuration, the lifting of degeneracy between the $Z_{3}$ vacua renders the string non-static. It will move towards the region of meta-stable states so as to reduce the free energy of the system [62]. In the presence of dynamical quarks, near $T_{c}$, only the $\theta=0$ state is observed in the deconfined phase. The states with $\theta=\pm 2\pi/3$ appear as metastable states only above a certain temperature $T_{m}$ [56]. This suggests that the string configuration will become unstable and ”melt” as the temperature drops below $T_{m}$. Analysis within the framework of the PNJL model at zero baryon chemical potential shows that if metastable states were to form in heavy-ion collision(HIC) experiments, they will survive until the temperature falls below $T_{m}$ [57]. This suggests that string configurations may potentially affect the dynamics of QGP above $T_{m}$ in HIC.
@@ II $Z_{3}$-string in pure $SU(3)$ gauge theory
In this section, we briefly discuss the $Z_{3}$ symmetry in pure $SU(3)$ gauge theory at finite temperature and how topological defects follow from the Polyakov loop effective potential. In pure $SU(3)$, the gauge fields $A_{\mu}$ at any point in space-time, are matrices in color space, i.e., $A_{\mu}=A^{a}_{\mu}T^{a}$, where $T^{a},a=1,2...8$, are generators of $SU(3)$ [65]. The partition function in the path-integral formulation is given by,
@@
Consequently, the effective potential has three degenerate minima, as illustrated in Fig.1. In the complex $L$-plane, the point $L=0$ corresponds to a saddle point of the effective potential, while the minima occur for $|L|\neq 0$ with the polar coordinate(phase) $\theta=0,2\pi/3$, and $2\pi/3$ [51, 43], which we denote by $L_{0},L_{1}$ and $L_{2}$ respectively. Given this, there are three possible domain walls, i.e., $L_{ij}$ interpolating $L_{i}$ and $L_{j}$. A static domain wall solution, on the yzyz-plane, is obtained by solving the following field equations,
${d^{2}l_{i}\over dx^{2}}={\partial V\over\partial l_{i}},\penalty 10000\ i=1,2,$ (6)
@@
$\lim_{x\to-\infty}L(x)=L_{i},\penalty 10000\ \lim_{x\to+\infty}L(x)=L_{j}.$
@@ III $SU(3)$ gauge theory on the lattice
We consider the standard Wilson action [28], as the discrete formulation of the action, Eq.2 , on a four dimensional Eucledian lattce, {align} S = β∑_n,μ¿ν Tr[1- 13 Re U_μν], where $U_{\mu\nu}=U_{\mu}(n)U_{\nu}(n+\hat{\mu}){U^{\dagger}}_{\mu}(n+\hat{\nu}){U^{\dagger}_{\nu}(n)}$ is the standard Wilson plaquette with $\mu,\nu=1,2,3,4$, $n=(\vec{n},n_{4})$. $\hat{\mu}$ is a vector of length equal to the lattice spacing($a$) in the $\mu$-th direction. $U_{\mu}(n)=\exp[igaA_{\mu}(n)]$ is the link variable, resides on the link connecting the lattice points $n$ and $n+\hat{\mu}$. $\beta=6/g^{2}$ is the lattice gauge coupling that regulates the lattice constant ($a$) and the temperature scale through the $\beta$-function.
@@
To compute the free energy of strings, appropriate boundary conditions are imposed on the lattice to induce a specific string configuration, as discussed in the next section. Specifically, we evaluate the free energy difference between the configurations with and without strings. The free energy can be calculated from the partition function, F=TlnZF=-TlnZ, where $Z$ is the partition function and $T$ is the temperature, but in practice, the direct evaluation of the partition function in lattice simulations is avoided as it is difficult. Instead, we evaluate the derivative of the free energy $w.r.t.$ the lattice gauge coupling $\beta$, which is related to the action difference between simulations with and without the string, i.e., {align} ∂∂β (FT) = 1β ⟨ΔS ⟩, where $\Delta S=S_{1}-S_{0}$, and $S_{1}$ and $S_{0}$ are actions with and without string respectively. The free energy is obtained by integrating the expectation value,
$\frac{F}{T}{\Big|}_{\beta_{c}}^{\beta}$ $=$ $\int_{\beta_{c}}^{\beta}d\beta\penalty 10000\ \left<{1\over\beta}\Delta S\right$> (7)
$=$ $\int_{\beta_{c}}^{\beta}d\beta\penalty 10000\ \left<\Delta\left(\sum_{n,\mu>\nu}Tr[1-\frac{1}{3}Re\penalty 10000\ U_{\mu\nu}]\right)\right$>